OpenStudy (anonymous):

(medal +fan :D and cookie) 2. -23= -2a-10+ a (1 point) a. a= 31 b. a= 13 c. a= -11 d. a = 5 please explain :)

OpenStudy (anonymous):

You can start by isolating the variable. Do this by moving the (-10) to the other side of the equal sign. Because (-10) is negative, you would add it to both sides. (+10) -23 = -2a - 10 (+10) + a This will give you -13 = -2a + a. Do you know where to go from here?

OpenStudy (anonymous):

erm... not really. :/ let me try something

OpenStudy (anonymous):

where did you get -10? that was suppose to be subtract

OpenStudy (anonymous):

You can think of it as subtracting or adding a negative. Whenever you are subtracting a number you are actually adding a negative number. So -10 is really +(-10). If that's confusing, you can also just think about what you can do to cancel a number out. If you are subtracting 10, what opposite operation can you perform to cancel that out? The opposite of subtraction is addition and whatever you do to one side of the equation you must do to the other side of the equation. So let's start over. -23 = -2a - 10 + a. You want to isolate the variable, which means getting the variable by itself on one side of the equation. This is how you find out what a is going to equal. To isolate the variable, you have to get rid of all numbers that don't contain a variable. In this equation, 10 does not have a variable so you need to get rid of it. Because you are subtracting 10, you have to now add it to cancel it out. -23 = -2a - 10 (+10) + a. What you do to one side of the equation you have to do to the other side of the equation, so you would have to add 10 to the other side. This will give you the equation (+10) - 23 = -2a - 10 (+10) + a. Does this make more sense?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

yes it makes sense so far

OpenStudy (anonymous):

Good :) So from here you would combine like terms. (+10) - 23 = -2a - 10 (+10) + a -13 = -2a + a You still have to get the variable by itself which means getting rid of all other numbers. You can always subtract or add like terms, including variables which are the same. So let's focus on combining the like terms of -2a + a. You can think of this a -2a + 1a. Simply, what is -2 + 1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

So -2a + a = -1a. Or -a. -13 = -a. To make this a positive number, you would simply change the signs. 13 = a. Or a = 13.

OpenStudy (anonymous):

a= 13 :)